OpenStudy (anonymous):

(5x-3)^2/3 =9 please help

OpenStudy (anonymous):

to make your life easier raise both side to power of three. try to solve the rest on your own or continue reading (5x - 3)^2 = 9^3 (5x - 3)^2 = 3^6 //square root of both side 5x - 3 = \[\pm\] 3^3 case 1: 5x -3 = 3^3 5x = 9 +3 x = 12/5 case 2: 5x -3 = - 3^3 5x = -9 +3 x = -6/5

OpenStudy (anonymous):

THANK YOU SO MUCH

OpenStudy (anonymous):

\[(5x-3)^{2}/3=9\]\[(5x-3)^{2}=3(9)\]\[(5x-3)=\pm \sqrt{27}\]\[5x=3\pm \sqrt{27}\]\[x=(3\pm \sqrt{27})/5\]

OpenStudy (anonymous):

my homework is sayint the anser is not correct

OpenStudy (anonymous):

mayb i can send u an attacment of how it look

OpenStudy (anonymous):

vivki do you mean (5x - 3)^(2/3) = 9 or ((5x - 3)^2)/3 = 9

OpenStudy (anonymous):

im trying to upload an attachment

OpenStudy (anonymous):

Just use equation

OpenStudy (anonymous):

how do u do fraction exponent on here

OpenStudy (anonymous):

Just hit the equation buton. Then look for the \[x^{a}\]

OpenStudy (anonymous):

Type x and then the x^a button.

OpenStudy (anonymous):

use brackets to show fraction and 4th button from tom shows you exponent

OpenStudy (anonymous):

\[solve (5x-3)_{3}^{2}=9\]

OpenStudy (anonymous):

and then it says enter the answer as x1,x2

OpenStudy (anonymous):

Do you know what the answer is exactly?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

just sent the attachment of exactly what the problem is

OpenStudy (anonymous):

OK then Safdari's answer should be correct.

OpenStudy (anonymous):

which answer?

OpenStudy (anonymous):

x=12/5 and x=-6/5

OpenStudy (anonymous):

i tried both and it still says incorrect idk y

OpenStudy (anonymous):

I didn't really look at the arthmetic. Give me a moment to check it.

OpenStudy (anonymous):

ok thank you so much for your time

OpenStudy (anonymous):

OK, Safdari had the method correct, but didn't do the arithmetic correctly. The answers should be x=6 and x=-24/5

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

mayb my homework is looking for something else im not sure. but thanks alot for your efforts

OpenStudy (anonymous):

marshall thanks for correcting, marshal is right but I think the right answer is case 1: 5x -3 = 3^3 5x = 27 +3 x = 30/5 x=6 case 2: 5x -3 = - 3^3 5x = -27 +3 x = -25/5 x = -5

OpenStudy (anonymous):

vicki try 6 and -5

OpenStudy (anonymous):

:( no luck

OpenStudy (anonymous):

-27+3=24

OpenStudy (anonymous):

Yikes! -27+3=-24

OpenStudy (anonymous):

are you guys familiar with educosoft math portal ? that is what im using

OpenStudy (anonymous):

there is no place for + sign

OpenStudy (anonymous):

Mother of careless I am I don't know what educosoft math is. and sorry for giving wrong answers over and over

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

Well we gave it a good try. Sorry we could be more helpful.

OpenStudy (anonymous):

mayb i can give u my login info so u can see for yourself how it is?

OpenStudy (anonymous):

oh ok no problem thank you very much

OpenStudy (anonymous):

I found something try this -24/5 = -4.8 try -4.8 and 6

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

no

OpenStudy (anonymous):

:( I have no idea

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

i do have other math homework i need help on. know of anywhere online i can pay for help?

OpenStudy (anonymous):

if I were you, I would just use MIT open course and learn them then I would do all homework myself. Send your homework here, maybe I can do something with it, this time. Trust me learning is more important, and asking some one to solve your assignment will not help you in future

OpenStudy (anonymous):

go to www.educosoft.com my login info is camille590@yahoo.com pw: 29870v

OpenStudy (anonymous):

what should I do? where are the questions?

OpenStudy (anonymous):

click assesments on the left

OpenStudy (anonymous):

and then u will see all of the ones with no score that says "take" or resume

OpenStudy (anonymous):

do you see???

OpenStudy (anonymous):

Yes, I have to answer the red ones?

OpenStudy (anonymous):

red what?

OpenStudy (anonymous):

red questions ?

OpenStudy (anonymous):

yes if they are gree they are correct red means wrong

OpenStudy (anonymous):

if my laptop powers of il get back on within a few mins just a head up . it overheats sometimes lol

OpenStudy (anonymous):

Okay, which of them do I have to do?

OpenStudy (anonymous):

any one u want they all need to be done really but which ever ones u feel up for..

OpenStudy (anonymous):

any luck?

OpenStudy (anonymous):

sorry I screwed one of the quiz :(

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

I got 84 I guess I stop here

OpenStudy (anonymous):

ooo u scared me 84 is ok with me

OpenStudy (anonymous):

THANKS ALOT :)